### v1

# Tuning Language Models for Robust Prediction of Diverse User Behaviors

Abstract
Predicting user behavior is essential for intelligent assistant services, yet deep learning models often struggle to capture long-tailed behaviors. Large language models (LLMs), with their pretraining on vast corpora containing rich behavioral knowledge, offer promise. However, existing fine-tuning approaches tend to overfit to frequent “anchor” behaviors, reducing their ability to predict less common “tail” behaviors. In this paper, we introduce BehaviorLM, a progressive fine-tuning approach that addresses this issue. In the first stage, LLMs are fine-tuned on anchor behaviors while preserving general behavioral knowledge. In the second stage, fine-tuning uses a balanced subset of all behaviors based on sample difficulty to improve tail behavior predictions without sacrificing anchor performance. Experimental results on two real-world datasets demonstrate that BehaviorLM robustly predicts both anchor and tail behaviors and effectively leverages LLM behavioral knowledge to master tail behavior prediction with few-shot examples.
1 Introduction
Most people lead routine lives shaped by habits, but also display short-term bursts of context-driven activities. Accurately predicting users’ next behavior is essential for intelligent assistant services across web platforms and smart devices Chung and Lee (2018); Tulshan and Dhage (2019); Savcisens et al. (2023); Rychalska et al. (2023).
With the accumulation of user behavior data, deep learning has replaced rule-based methods Zhang and Dai (2018); Zhang et al. (2019); Li et al. (2022), with transformer-based models excelling at capturing sequential patterns Kang and McAuley (2018); Pu et al. (2018); Sun et al. (2019); Savcisens et al. (2023). However, their reliance on extensive training data leads to high costs. In this context, Large language models (LLMs) Brown (2020); Zhao et al. (2023) offer a promising alternative by leveraging pretrained knowledge of human behavior, showing potential in interpreting user intent and generating accurate predictions Wu et al. (2024); Zhao et al. (2024).
Recent works adapt LLMs for behavior prediction by converting behavioral sequences into text and fine-tuning them to predict user behaviors Li et al. (2023); Bao et al. (2023); Liao et al. (2024); Kim et al. (2024). Early methods Bao et al. (2023) combined general instruction tuning with lightweight behavior fine-tuning, and subsequent research Liao et al. (2024) enhanced LLMs with embeddings from traditional models. While expressing user behaviors as text appears intuitive Li et al. (2023), the long-tailed distribution of behaviors presents a significant challenge Liu et al. (2024): a user’s daily life is predominantly characterized by a small subset of frequent behaviors serving as “anchors”, while other behaviors occur far less commonly. Recent work Liu et al. (2024) explores hybrid methods, combining LLMs’ profiling capabilities with traditional models to address this issue. However, capturing long-tailed distributions remains a persistent challenge for LLMs. Our empirical observations reveal that while an 8B-parameter LLM fine-tuned on behavior data outperforms untuned GPT-4 on anchor behaviors, it still underperforms on tail behaviors. This highlights a fundamental limitation: fine-tuning alone does not uniformly enhance prediction capabilities across diverse behaviors.
In this paper, we empirically investigate LLMs’ uneven performance across behavior types and make a surprising discovery: an LLM fine-tuned only on anchor behaviors can still generalize well to tail behaviors in a zero-shot manner, even better than full data fine-tuning. Motivated by this finding, we propose BehaviorLM, a two-stage fine-tuning progressive framework designed to better model long-tailed user behaviors. In the first stage, we fine-tune the LLM using anchor behaviors and a general conversational corpus in a multi-task setup to preserve general behavioral knowledge and prevent bias. In the second stage, we fine-tune the model on a balanced subset of all behaviors, selected by sample difficulty, to improve tail behavior predictions without compromising anchor performance. Unlike existing methods that rely on auxiliary data such as embeddings from traditional deep learning-based models, increasing training complexity and limiting generalization, our approach leverages the intrinsic semantic knowledge of LLMs to overcome these challenges.
Our main contributions are as follows:
This paper introduce a novel progressive fine-tuning method for adapting LLM into the robust predictor of user daily behaviors, addressing the challenge of long-tailed user preferences.
We propose a progressive fine-tuning method that effectively leverages LLMs’ general behavioral knowledge for accurate and efficient behavior prediction.
Extensive experiments on two real-world datasets show BehaviorLM achieves up to 27.4%/20.4% accuracy improvements on tail behaviors, offering 100 higher sample efficiency than traditional transformer models and enabling few-shot prediction with as few as 20 examples. Ablation studies further validate the design of our method.
2 Motivation
We first give a formal definition of our research problem and then present our novel observations that motivate the methodology design of BehaviorLM.
2.1 Problem Formulation
Next behavior prediction refers to the task of predicting a user’s next behavior, , given a chronologically ordered sequence of their most recent historical behavioral events. Each event indicates that a specific behavioral event occurred at location and time . The behavioral event specifically refers to daily activities—such as exercise or gaming—rather than fine-grained actions like picking up a cup. The location indicates a semantic indicator such as "home" or "workplace." The time includes both date and hour information.
Given a collected dataset , our goal is to train a prediction model capable of predicting the next behavioral event, i.e.,
2.2 Investigating LLM Fine-tuning Performance on Diverse User Behavior Dataset
In this subsection, we evaluate the behavior prediction performance of fine-tuning approaches for LLMs using a real-world dataset of 37 smartphone user behaviors (detailed in Section 4). As shown in Figure 1(a), there is a significant class-imbalance issue, with behaviors divided into anchor (occurrence >1%) and tail categories. Despite only 16 anchor behaviors, they represent over 97% of the data, with an average anchor-to-tail occurrence ratio of 42.44 (0.97/16 vs. 0.03/21) and a maximum exceeding 2,500 (43.6% vs. 0.0163%).
To further analyze the semantic meaning and similarity of diverse behaviors from the LLM’s perspective, we use the pretrained Llama-8B v3.1 Dubey et al. (2024) to generate behavior embeddings and reduce them from 4096 to 2 dimensions using PCA Abdi and Williams (2010). As shown in Figure 1(b), anchor behaviors serve as semantic anchors in the latent space, with tail behaviors clustering around them. For example, watching videos, especially short ones, is highly frequent, while nearby behaviors like watching sports or animation occur less often and appeal to fewer users. Similarly, public transportation is a more common commuting behavior than hailing a taxi.
To investigate LLM behavior prediction performance, we fine-tune Llama-8B v3.1 on the previously described behavior dataset and compare average prediction accuracy for anchor and tail behaviors using an equal number of samples per behavior type (Figure 1(c)). For reference, GPT-4 (gpt-4o-2025-03-26) achieves accuracy scores of 0.45 for anchor behaviors and 0.33 for tail behaviors. In its base form, Llama-8B v3.1 performs worse than GPT-4 for both anchor and tail behaviors. While fine-tuning improves its performance on anchor behaviors, surpassing GPT-4, it continues to underperform on tail behaviors. GPT-4’s strong overall performance, achieved without behavior-specific fine-tuning, suggests current fine-tuning methods overly favor anchor behaviors at the expense of tail behaviors, likely due to the severe class imbalance shown in Figure 1(a).
Motivated by these observations, we explored whether addressing class imbalance during fine-tuning could better utilize the LLM’s inherent understanding of user behaviors. We fine-tuned the model exclusively on anchor behavior data, omitting tail behaviors. Notably, this approach preserved strong performance on anchor behaviors while enabling robust zero-shot prediction for tail behaviors, achieving an accuracy of 0.39—significantly higher than the 0.29 accuracy from traditional fine-tuning. This suggests that focusing on anchor behaviors allows the LLM to capture fundamental user behavior patterns and leverage its general knowledge to generalize to semantically related, unseen behaviors. This insight underpins our progressive fine-tuning approach for BehaviorLM.
3 Method
3.1 The BehaviorLM Framework
3.1.1 Behavior Prediction as Language Modeling
To frame the next behavior prediction task as a language modeling problem, we design a specific prompt and apply instruction fine-tuning to the LLM.
Prompt design. We employ a text-only prompt to convert user behavior history and the next behavior (if provided) into text-based metadata suitable for instruction tuning. The prompt consists of five components:
1.
Task definition: A concise description of the task, e.g., "Predict the user’s next behavior from the candidates.
2.
Role-playing instruction: Defines the LLM’s role and task Shanahan et al. (2023), e.g., "You are a smart mobile assistant that can infer the user’s behavior preferences based on their history."
3.
Historical behavior sequence (Input): Transforms each element in into a textual description, e.g., "(1, 16, home, Exercise), (1, 20, home, Gaming), (1, 20, home, Video)."
4.
Candidate set (Input): Specifies all possible behaviors for prediction, e.g., "Weather Check, Music, Cycling, Reading."
5.
Next behavior (Output): The ground truth label for the LLM to predict, e.g., Gaming.
This design transforms user behavior data into a format optimized for LLM instruction fine-tuning.
Instruction Fine-Tuning LLMs. To quickly adapt an LLM for behavior prediction tasks, we use instruction fine-tuning Ouyang et al. (2022). The user behavior dataset is converted into a text-based instruction dataset , where and are natural language representations of inputs and outputs. The LLM is optimized with the next-token prediction loss:
where are the LLM parameters, is the -th token of , and represents tokens preceding . Moreover, we adopt a well-known parameter-efficient fine-tuning technique, i.e., LoRA Hu et al. (2021), to fine-tune the LLM more efficiently.
3.1.2 Progressive Fine-Tuning with Behavior Data
As illustrated in Figure 1(a), user behavior data often exhibits a significant class imbalance, with many behaviors occurring infrequently, leading to a long-tailed distribution. When the task of predicting a user’s next behavior is reframed as generating the corresponding language token, this imbalance persists, causing poor performance on infrequent (tail) behaviors, as shown by the results of "Llama3.1-8B-FT" in Figure 1(b). Interestingly, we observe that fine-tuning an LLM on a subset of frequently occurring behaviors (anchor behaviors) enables it to act as a surprisingly effective zero-shot predictor for less frequent (tail) behaviors, often outperforming fine-tuning on the full dataset. Inspired by this, we propose a progressive fine-tuning strategy to enhance BehaviorLM’s ability to predict a diverse range of user behaviors.
The proposed fine-tuning strategy consists of two progressive stages (Figure 2). In the first stage, referred to as A-Tuning, the LLM is fine-tuned using anchor behavior data to specialize in user behavior prediction while retaining its inherent rich knowledge of long-tailed behaviors. In the second stage, referred to as B-Tuning, the model is further fine-tuned on a class-balanced dataset covering all behaviors in a few-shot manner, enhancing its predictive capability for tail behaviors. The first stage helps the LLM become a specialist in predicting anchor behaviors, while the second stage transforms it into a generalist capable of predicting both anchor and tail behaviors.
3.2 Multitask-Enhanced Anchor Behavior Tuning (A-Tuning)
We divide the instruction fine-tuning data, , into two subsets: , containing labels for anchor behaviors, and , encompassing the rest. In this stage, the LLM is fine-tuned exclusively on . As illustrated in Figure 1, anchor behaviors reflect core patterns of a user’s daily life. Therefore, fine-tuning on anchor behaviors provides the LLM with a foundational understanding of these underlying patterns, enabling accurate next-anchor behavior predictions. Moreover, since the LLM already captures general behavioral knowledge from pretraining, fine-tuning on anchor behaviors facilitates generalization to unseen but semantically related behaviors. Additionally, compared to fine-tuning on the full dataset , this approach mitigates bias toward overrepresented behavior types.
To further improve generalization from anchor to tail behaviors, we proposed a multi-task fine-tuning approach. Beyond learning to predict next behaviors in , we maintain the model’s general task-solving ability, which our experiments later confirmed to be helpful for behavior prediction tasks. To achieve this, we jointly fine-tune the model on an auxiliary dataset , comprising daily user-ChatGPT conversations Zheng et al. (2023). This integrates a conversation generation task alongside the primary behavior prediction task. We control the auxiliary task’s influence by adjusting the size of relative to , using a ratio . Additionally, in practice, we filter out excessively long conversations from to ensure that the prompts for both tasks are comparable in length.
3.3 Difficulty-based Data Selection for Balanced Behavior Tuning (B-Tuning)
In the second stage, we reintroduce the tail behaviors and combine them with to create a class-balanced fine-tuning dataset. Since the LLM fine-tuned during the A-Tuning stage already serves as a good zero-shot predictor for tail behaviors, we believe that a small amount of instruction fine-tuning data covering all behavior types should suffice to build a robust user behavior predictor. However, to achieve this, the quality and informativeness of the selected samples play a crucial role. Specifically, we search for difficult samples from the following two dimensions for effectively fine-tuning the LLM in a few-shot way.
Obtaining Mispredicted Samples. One simple way to measure sample difficulty is by scoring samples with an intermediate model, and those with the wrong predicted labels are more difficult than those correctly predicted Bengio et al. (2009). In our approach, we utilize the LLM fine-tuned during A-Tuning as the scorer and extract mispredicted samples from and for each behavior type.
Contrastive Data Selection. Inspired by contrastive learning Liu et al. (2021), we identify challenging samples by evaluating the distinguishability between their predicted and ground truth labels. Specifically, we select mispredicted samples with lower distinguishability scores , computed as follows:
where denotes the predicted behavior, denotes the ground truth label, and denotes behavior category of , i.e., anchor or tail. For each type of behavior, we only select those mispredicted samples with . This follows the idea that anchor behaviors (tail behaviors) are much more difficult to discriminate from themselves than their counterparts.
Finally, we combine the above two dimensions to measure sample difficulty and select a fixed number of samples for each type of behavior to create the class-balanced fine-tuning dataset§§§Since is usually small, we do not have to score each sample with the LLM to obtain mispredicted samples, which can be time-costly.. By adopting the difficulty-based data selection strategy, we can prompt the LLM to capture more refined behavior knowledge in a sample-efficient manner and significantly enhance its prediction accuracy across tail behaviors without compromising those of anchor behaviors.
We also summarize the designed progressive fine-tuning strategy for BehaviorLM in Algorithm 1.
4 Experiment
4.1 Experiment Settings
4.1.1 Datasets
We evaluated our model on two real-world user behavior datasets: App usage dataset Li et al. (2022): This open-source dataset records 1,753 users’ interactions across various apps over one week. Grouped by app functions, it includes 24 behaviors reflecting key daily activities. Behavior dataset: This large-scale dataset is sourced from mobile device §§§https://www.honor.com/global/usage logs, capturing user interactions and their corresponding behavioral intentions. After processing, it covers 37 daily behaviors across various scenarios, including learning, work, entertainment, and leisure.
More details about the datasets and preprocessing can be seen in Appendix 3.
4.1.2 Evaluation Protocols and Metrics
To thoroughly evaluate our model, we adopt two evaluation protocols. First, for next behavior prediction Bao et al. (2023); Liao et al. (2024), we split users 8:1:1 into training, validation, and testing sets to assess overall performance. Second, to evaluate tail behaviors, we follow the common practice of long-tailed learning Liu et al. (2019); Shi et al. (2024), construct a balanced test set with 500 samples per behavior, ensuring fair evaluation of tail behaviors. Note this test set is derived from the original Behavior and App datasets, independent of the first protocol. All experiments are repeated five times, with results averaged.
Overall, we adopt six commonly used metrics. For next behavior prediction, we adopt weighted precision () and weighted recall (), accounting for the overall performance by considering the proportion of behavior categories. Note that the weighted recall () used here corresponds to the common recommendation metric Hit-Rate. For long-tailed learning, following the settings of Liu et al. (2019); Shi et al. (2024), we evaluate four accuracy metrics based on behavior occurrence frequencies: category-average accuracy across all behaviors (), head-category accuracy for behaviors with >5.0% frequency (), medium-category accuracy for those between 1.0% and 5.0% frequency (), and tail-category accuracy for the remaining low-frequency behaviors (). These six metrics assess the model’s robustness across diverse behaviors and its ability to handle imbalanced datasets effectively. Detailed calculation formulas for the six metrics are provided in Appendix B.
4.1.3 Baselines
We selected the following seven representative algorithms to compare with our proposed algorithm, covering traditional methods (SASRec Kang and McAuley (2018) and Bert4Rec Sun et al. (2019)), LLM-enhanced methods (PITuning Gong et al. (2024), LLM-ESR Liu et al. (2024) and AlphaFuse Hu et al. (2025)), LLM-based methods (GPT4o Achiam et al. (2023), A-LLMRec Kim et al. (2024), TALLRec Bao et al. (2023) and LLaRa Liao et al. (2024)). The details of baselines are provided in Appendix C.
4.1.4 Implementation Details
We selected LLama3.1-8B Dubey et al. (2024) as the backbone for our experiments. Our experiments utilized the AdamW optimizer with a cosine annealing learning rate schedule. And we conducted LoRA fine-tuning and parallel training acceleration using the open-source LLM instruction fine-tuning library, llama-factory Zheng et al. (2024). More details about the implementation are in the Appendix D.
App Usage dataset | Behavior Dataset | ||||||||||||
Category | Model | ||||||||||||
Traditional | SASRec | 0.5309 | 0.5759 | 0.2752 | 0.5255 | 0.2567 | 0.1733 | 0.4818 | 0.5507 | 0.1531 | 0.4083 | 0.2159 | 0.0941 |
Bert4Rec | 0.3452 | 0.5400 | 0.0962 | 0.5290 | - | - | 0.1908 | 0.4368 | 0.0290 | 0.2500 | - | - | |
LLM-Enhanced | PITuning | 0.5837 | 0.5133 | 0.2910 | 0.5029 | 0.3721 | 0.1066 | 0.5829 | 0.5057 | 0.2121 | 0.5030 | 0.3545 | 0.0522 |
LLM-ESR | 0.5437 | 0.5906 | 0.2779 | 0.5615 | 0.2467 | 0.1750 | 0.5325 | 0.5750 | 0.1633 | 0.4240 | 0.1772 | 0.0977 | |
AlphaFuse | 0.5621 | 0.6024 | 0.2873 | 0.5706 | 0.2552 | 0.1881 | 0.5428 | 0.5857 | 0.1729 | 0.4335 | 0.1874 | 0.1093 | |
LLM-Based | LLama-NT | 0.5467 | 0.5346 | 0.3736 | 0.5335 | 0.3685 | 0.3000 | 0.5091 | 0.4226 | 0.2433 | 0.4142 | 0.2467 | 0.2040 |
GPT4o | 0.5872 | 0.5678 | 0.4557 | 0.5410 | 0.4642 | 0.4025 | 0.5735 | 0.5660 | 0.3561 | 0.4575 | 0.3703 | 0.3248 | |
A-LLMRec | 0.5908 | 0.6154 | 0.3514 | 0.5815 | 0.3539 | 0.2333 | 0.5545 | 0.5856 | 0.2526 | 0.4567 | 0.3267 | 0.1601 | |
LLaRA | 0.6074 | 0.6256 | 0.4455 | 0.5970 | 0.4468 | 0.3683 | 0.5892 | 0.6099 | 0.3458 | 0.5175 | 0.3661 | 0.2932 | |
TALLRec | 0.6173 | 0.6306 | 0.4397 | 0.6205 | 0.4328 | 0.3580 | 0.5938 | 0.6141 | 0.3425 | 0.5175 | 0.3636 | 0.2908 | |
BehaviorLM | 0.6343 | 0.6364 | 0.5379 | 0.6220 | 0.5240 | 0.5125 | 0.6145 | 0.6261 | 0.4238 | 0.5392 | 0.4353 | 0.3912 | |
Improv. | 2.8% | 0.9% | 18.0% | 0.2% | 12.9% | 27.4% | 3.5% | 2.0% | 19.0% | 4.2% | 17.6% | 20.4% |
4.2 Overall Performance
We compare the performance of BehaviorLM with other baseline methods across two evaluation settings, and the results are summarized in Table 1. Our key observations are as follows:
BehaviorLM consistently outperforms across all metrics on both datasets. Notably, under the long-tailed learning evaluation protocol, BehaviorLM improves Overall Accuracy by up to 18.0% on the App dataset and 19.0% on the Behavior dataset. This performance is driven by its superior ability to predict tail behaviors, surpassing the best baseline by 27.4% (App) and 20.4% (Behavior).
Fine-tuning LLMs for behavior prediction requires addressing the severe long-tailed distribution. Existing LLM fine-tuning methods, such as LLaRA Liao et al. (2024) and TALLRec Bao et al. (2023), which directly fine-tune LLMs on all behavior types, fail to outperform non-tuned GPT4o on Medium and Tail categories. Similarly, hybrid LLM-enhanced methods that integrate LLMs into traditional models lack robustness across both anchor and tail behaviors. BehaviorLM overcomes this challenge through a novel progressive tuning strategy, first mastering anchor behaviors and then generalizing to tail behaviors.
Behavioral knowledge embedded in LLMs significantly enhances behavior prediction. Traditional deep learning-based solutions, such as SASRec Kang and McAuley (2018) and Bert4Rec Sun et al. (2019), struggle to compete with LLM-based prediction models because they cannot leverage this knowledge and perform poorly when fine-tuning data is limited. Additionally, Bert4Rec performs particularly poorly under long-tailed evaluation settings, failing to make accurate predictions on less frequently occurring behaviors (Medium and Tail categories).
4.3 Investigating the Effect of Behavioral Knowledge
Our evaluation in the previous subsection highlights the significant performance improvement brought by the LLM’s behavioral knowledge, as evidenced by the superior performance of LLM-based methods over traditional deep learning approaches. In this subsection, we conduct a more detailed analysis from this perspective. It is well-established that the more parameters an LLM contains, the greater its capacity to store knowledge learned from its pretraining corpus. Therefore, we examine the impact of model size on the LLM’s predictive capability by fine-tuning three versions of BehaviorLM with different backbones: Qwen-1.5B-v2, Llama-8B-v3.1, and Llama-70B-v3.1. Here, the proportion of auxiliary task data is controlled at 5% for all models.
4.3.1 Performance robustness
We investigate whether this behavioral knowledge enables robust prediction across both anchor and tail behaviors, as well as under diverse few-shot settings.
In Figure 3(a), We evaluate BehaviorLM on anchor and tail behaviors (Behavior dataset) across model sizes and tuning strategies. Scaling from 1.5B to 70B significantly improves accuracy—for anchors (0.415 → 0.451) and tails (0.354 → 0.385)—demonstrating robustness. However, without our progressive tuning (A-Tuning + B-Tuning), larger models yield only marginal gains (blue curve in Figure 3(a)), highlighting the effectiveness of our approach in leveraging LLM behavioral knowledge.
In Figure 3(b), we vary the number of few-shot examples used in the B-tuning stage for BehaviorLM-8B and plot the performance curve. It is evident that with fewer than 20 examples per behavior type, BehaviorLM quickly learns to make robust predictions, demonstrating its ability to grasp behavior patterns efficiently even in low-data settings.
Additionally, we conduct efficiency experiments demonstrating that the behavioral knowledge in LLMs offers over 100× greater sample efficiency than traditional transformers. Experimental details are in the E.
4.4 Ablation Study
In Table 5, we evaluate the contribution of each design component to overall performance through ablation studies on the Behavior dataset. Specifically, we examine the performance drop when: (1) removing the auxiliary task in A-tuning, (2) replacing difficulty-based data selection with uniform selection, (3) performing non-contrastive data selection by setting the distinguishability score , and (4)replacing A-tuning with fine-tuning on all behaviors, as done in previous works Bao et al. (2023); Liao et al. (2024). The ablation experiment results of the APP dataset are detailed in the Appendix F.
Key findings are as follows: All design components enhance prediction performance. For head behaviors, B-tuning’s data selection is most impactful by improving behavior differentiation. For medium and tail behaviors, A-tuning is essential. Without it, accuracy drops sharply, as direct fine-tuning dilutes performance on tail behaviors.
Dataset | BehaviorLM Variant | Overall | Head | Medium | Tail |
Behavior | A-Tuning w/o Aux. Task | -4.71% | -1.07% | -5.99% | -4.58% |
B-Tuning w/o DDS | -6.23% | -4.34% | -7.66% | -5.62% | |
B-Tuning w/o CDS | -5.50% | -3.24% | -6.83% | -5.02% | |
w/o A-Tuning | -10.00% | -2.77% | -11.74% | -10.61% |
4.5 Investigating the Necessity of Anchor Behavior Tuning
We have demonstrated the importance of progressive tuning for LLM-based behavior prediction. However, the necessity of first tuning on anchor behaviors remains unclear. To address this, we replace the A-tuning stage with a tuning stage focused solely on tail behaviors(here, we also include medium-frequency behaviors, as tail behaviors are relatively rare). As illustrated in Figure 4, this alternative approach leads to poorer prediction performance in both anchor and tail behaviors. This suggests that, since anchor behaviors represent the core structure of human daily life, prompting the LLM to follow a curriculum from anchor behaviors to tail behaviors is more effective than the reverse.
5 related works
5.1 LLM-Enhanced Behavior Prediction Models
User behavior prediction based on the most recent events is similar to sequential recommendation. The key difference is that behavior prediction focuses on recurring daily actions, while item recommendation emphasizes novel content. Due to limited research in behavior prediction, we draw on related work in sequential recommendation. Recent studies in recommendation explore knowledge alignment between language and recommendation domains. For example, A-LLMRec used multi-task training for embedding alignment Kim et al. (2024), PLM-Rec applied mutual information maximizationGeng et al. (2022) , PITuning introduced PITuning for cross-modal pattern extractionGong et al. (2024), and LLM-ESR proposed a cross-attention mechanism for sequence alignmentLiu et al. (2024). While these methods enhance traditional models by aligning language and recommendation knowledge, they underutilize LLMs’ zero-shot and few-shot generalization. Our approach addresses this gap with progressive fine-tuning, preserving general behavioral knowledge while improving prediction of infrequent long-tail behaviors without sacrificing performance on frequent ones.
5.2 LLM-Based Behavior Prediction Models
Traditional behavior prediction models often rely on embedding-based neural networks, which lack interpretability. LLMs, with their rich world knowledge and reasoning capabilities, present a more interpretable alternative for user behavior prediction. Tallrec first showcased LLMs’ strong few-shot performance in recommendation tasks Bao et al. (2023), while Llara improved fine-tuning through curriculum learning, progressing from simpler to more complex tasks Liao et al. (2024). Similarly, Recexplainer aligned LLMs with traditional recommendation models for user-specific, customizable explanations Lei et al. (2024) . Other notable frameworks include the Chain-of-Planned Behavior Shao et al. (2024), which captured spatial-temporal dynamics, and the Transition paradigm TransRec Lin et al. (2024), combining multiple identifiers to enhance LLM-based recommendations. However, these approaches mainly convert behavior sequences into text for LLM training, often overlooking critical distinctions between anchor and tail behaviors, which hampers their zero-shot generalization. Our method, BehaviorLM, directly addresses the long-tailed behavior distribution by employing progressive fine-tuning. This two-stage process significantly improves the prediction of tail behaviors without compromising anchor behavior performance.
6 Conclusion
In this paper, we leverage the rich behavioral knowledge in LLMs to tackle user behavior prediction, with a focus on long-tail behavior prediction. We propose a progressive tuning approach, where the LLM first learns frequent anchor behaviors before generalizing to rarer tail behaviors. Experiments on two real-world datasets show that BehaviorLM outperforms state-of-the-art methods, achieving up to 27%/20% improvement in long-tail behavior prediction, addressing a traditionally challenging aspect of behavior modeling. Analysis highlights that addressing the long-tailed behavior distribution is essential for effectively utilizing LLMs’ behavioral knowledge in fine-tuning.
Appendix A Details of Used Datasets
Behavior dataset: This large-scale dataset is derived from mobile phone usage logs. When users interact with their mobile phones, various types of logs are generated, desensitized, and reported with user consent. We select 37 daily behaviors that are reliably extracted from raw logs and also cover broad life scenarios, including activities related to learning, work, entertainment, leisure, etc. The dataset spans from March 1, 2024, to April 29, 2024, and consists of over 50 million behavior events from 24,133 anonymous users. We preprocess the dataset and construct samples in the format of “(last-20 events, next event)”. Since our target is fine-tuning the LLM instead of training from scratch, we only randomly select a subset (200,000) of data for experiments.
App usage dataset [16]: This dataset is an open-source resource that captures 1753 user interactions across various apps within one week. Given the large number of apps and the overlap in functionality among many of them, we processed the dataset further by merging apps with similar purposes. For example, Douyin and Kuaishou were grouped into a "watching videos" category. After similar preprocessing as the Behavior dataset, this dataset consists of 71 behavior events belonging to 24 categories.
Behavior | Proportion |
Watching video clip | 0.435981 |
Shopping online | 0.112217 |
Editing video | 0.05937 |
Reading books | 0.055293 |
Gaming | 0.044202 |
Watching video | 0.041266 |
Reading news | 0.036046 |
Taking photos | 0.035068 |
Taking subway | 0.031969 |
Payment | 0.029848 |
Listening Music | 0.02365 |
Checking Weather | 0.016963 |
Studying for course | 0.01419 |
Ordering takeout | 0.012722 |
Listening podcast | 0.012396 |
Conducting job recruitment | 0.010928 |
Taking notes | 0.003915 |
Checking calendar | 0.003591 |
Hailing taxi | 0.003262 |
Picking up the express delivery | 0.00261 |
Watching Match | 0.00213 |
Exercising | 0.002003 |
Singing | 0.001957 |
Watching Live-streaming | 0.001794 |
Cycling | 0.001794 |
Meeting online | 0.000816 |
Studying for driving test | 0.000652 |
Translating | 0.000652 |
Rentaling house | 0.000652 |
Watching animation | 0.000489 |
Booking ticket | 0.000489 |
Commuting check-in | 0.000323 |
Cooking | 0.000323 |
Drinking coffee | 0.000283 |
Rentaling car | 0.000163 |
The characteristics of both two datasets are presented in Table 4.
Dataset | # User | # Behavior Type | # Sample |
Behavior Dataset | 24,133 | 37 | 200,000 |
App Usage Dataset | 1,753 | 24 | 50,000 |
Appendix B Details of Used Metrics
The formula for :
The formula for :
The formula for :
The formula for :
The formula for :
The formula for :
Where represents the total number of classes, represents the total number of classes belonging to the head category,
Where represents the total number of classes belonging to the medium category,
Where represents the total number of classes belonging to the tail category.
True Positives denotes the number of samples correctly classified as class , False Positives represents the number of samples incorrectly classified as class , and False Negatives stands for the number of samples incorrectly classified as other classes instead of class . And and respectively refer to the precision and recall of class .
Appendix C Baselines
SASRec [12]. uses self-attention mechanisms to model user behavior sequences. It captures both short-term and long-term dependencies in sequential data, allowing it to focus on the most relevant items in the user’s interaction history for recommendation.
BERT4Rec [30]. models user behavior sequences using deep bidirectional self-attention. By jointly considering the context before and after an item, it predicts the randomly masked items within the sequence, achieving excellent predictive performance.
LLaRa [17] introduces a hybrid prompting method that integrates both world knowledge and behavioral patterns into item representations. It conducts curriculum prompt tuning to achieve modality alignment.
A-LLMRec [13] bridges the knowledge between the language and recommendation domains by training an alignment network with a variety of tasks, targeting both warm and cold-start scenarios.
PITuning [9] loads pre-trained Large Language Model (LLM) parameters to acquire textual knowledge and then designs an adaptive unlearning strategy to address the long-tail preference issue, achieving excellent performance in user behavior prediction.
LLM-ESR [19] leverages the semantic information from LLMs, proposes a dual-view modeling framework enhanced through embedding techniques to capture the nuances of long-tail items better, demonstrating strong performance across multiple datasets..
TALLRec [3] is one of the earlier methods to integrate Large Language Models (LLMs) with the recommendation domain. It employs a two-stage tuning process—Alpaca Tuning and Rec-Tuning—to finetune LLMs for recommendations, enabling effective and efficient adaptation of LLMs with only a small number of tuning samples.
AlphaFuse [11] is a simple yet effective language-guided learning strategy that addresses long-tail intent modeling by learning ID embeddings within the null space of language embeddings.
Appendix D Implementation Details
The hardware used in this experiment consists of 8 NVIDIA A100 40G GPUs. We selected LLama3.1-8B [7] as the backbone for our experiments. To ensure flexibility in model testing, we designed three distinct instruction formats, which are randomly sampled during both training and testing. Our experiments utilized the AdamW optimizer with a cosine annealing learning rate schedule, setting the warm-up proportion to 0.1. The maximum learning rate for cosine annealing was set to 1e-4, while both the minimum and initial warm-up learning rates were set to 1e-6. We conducted LoRA fine-tuning and parallel training acceleration using the open-source LLM instruction fine-tuning library, llama-factory [38]. All experiments were performed with a maximum of 8 training epochs and a batch size of 8, selecting the best-performing model on the validation set for testing. Detailed formats of the three instruction types are as followed:
Instruct1: This user has done behaviors [HistoryHere] in the previous. Day of the week, the hour, and the place of the next behavior are [next-intent-info], respectively. Choose the answer from the following behavior candidate set: [CansHere]. The answer is [Output].
Instruct2: The user’s historical behavior information sequence is: [HistoryHere]. Day of the week, the hour, and the place of the next behavior are [next-intent-info], respectively. Given the following behavior candidate set: [CansHere], recommend one intention for this user to do next. The intent you recommend is [Output].
Instruct3: The behavior history of this user is: [HistoryHere].Day of the week, the hour, and the place of the next behavior are [next-intent-info, respectively. Recommend a next intention for this user to do from the following behavior candidate set: [CansHere].The recommendation is [Output].
Appendix E Sample efficiency.
One significant advantage of leveraging the LLM’s general behavioral knowledge is that it reduces the need for fine-tuning on large-scale user behavior data, demonstrating strong sample efficiency. To validate this, we compare BehaviorLM-1.5B and BehaviorLM-8B with another transformer-based model trained from scratch, using the objective function from SASRec [12] and the model architecture from GPT2 [24] (the small version with 12 layers and 768 latent dimensions). Since the original Behavior dataset contains over 50 million events, we vary the sample size to observe its impact on prediction performance. As shown in Figure 5, BehaviorLM demonstrates a significant improvement in sample efficiency. The transformer-based model trained from scratch only outperforms BehaviorLM when trained on nearly all 50 million samples, while BehaviorLM is fine-tuned on just 200,000 samples—over two orders of magnitude fewer. This highlights the remarkable sample efficiency advantage provided by the LLM’s preexisting behavioral knowledge. Using 8xA100 (40G) GPUs, training BehaviorLM-8B on Behavior Dataset with 200k samples takes about 6 hours, while on APP Usage dataset (50000 samples) it takes no more than 2 hours. In contrast, training a SASRec model on a 50M sample dataset requires approximately 72 hours. Overall, although BehaviorLM has a lower training speed for each sample, it has a much higher sample efficiency than SASRec and thus takes much less time to achieve a good prediction accuracy.
Appendix F Ablation Study
Dataset | BehaviorLM Variant | Overall | Head | Medium | Tail |
App Usage | A-Tuning w/o Aux. Task | -3.83% | -2.50% | -2.17% | -6.25% |
B-Tuning w/o DDS | -7.80% | -5.80% | -6.06% | -10.42% | |
B-Tuning w/o CDS | -3.87% | -1.29% | -1.91% | -8.47% | |
w/o A-Tuning | -7.95% | -0.97% | -5.64% | -14.40% | |
Behavior | A-Tuning w/o Aux. Task | -4.71% | -1.07% | -5.99% | -4.58% |
B-Tuning w/o DDS | -6.23% | -4.34% | -7.66% | -5.62% | |
B-Tuning w/o CDS | -5.50% | -3.24% | -6.83% | -5.02% | |
w/o A-Tuning | -10.00% | -2.77% | -11.74% | -10.61% |
Appendix G Details of the Hyperparameter Experiment on Auxiliary Task Ratio
Auxiliary Task Data Ratio | overall | head | medium | tail |
0% | 0.404 | 0.528 | 0.410 | 0.373 |
2.5% | 0.410 | 0.528 | 0.415 | 0.380 |
5% | 0.424 | 0.534 | 0.435 | 0.391 |
10% | 0.412 | 0.525 | 0.416 | 0.385 |
20% | 0.409 | 0.512 | 0.411 | 0.358 |
As evident from Table 6, an appropriate amount of auxiliary task data can significantly enhance the model’s performance. Insufficient auxiliary task data fails to deliver notable improvements, whereas an excessive amount can disrupt the model’s predictive capabilities.